# Hidden memory and stochastic fluctuations in science

### v1

@@ III.4 The fractional stochastic differential equation
Now, we explore the infinitesimal-time limit $\Delta t\to dt$. To incorporate our set of Ansatze (i)(iv), we consider the following fractional SDE, which describes the evolution of latent attention:
@@
The time-dependent drift function $\alpha(t)$ determines the shape of the average attention history curve $X(t)$ (see Eq. (11) below). The parameter $\beta>0$, referred to as the volatility parameter—borrowing terminology from financial engineering—controls the magnitude of random fluctuations in $X(t)$ over the interval $dt$, reflecting external random events as discussed earlier.
As is evident, the fractional SDE (8) inherently incorporates our Ansatze (iii) and (iv) by definition. In particular, Ansatz (iii) satisfies the property of nonlinear preferential attachment. Compared to the SDE in the Black–Scholes model [51, *Merton73], this fractional SDE differs in two key aspects: Brownian motion is replaced by fractional Brownian motion, and the drift term $\alpha(t)$ is a function of time rather than a constant. If $\alpha$ were constant, this model would coincide with the fractional Black–Scholes model [66]. In addition, compared to the SDE considered in [23, Eqs. (14, S3)], this fractional SDE differs in that Brownian motion is extended to fractional Brownian motion, while the volatility remains a time-independent constant.
[…]
@@
Using the properties of fractional Brownian motion given by Eqs. (4, 5), we obtain $\mathbb{E}[R(t)]=A(t)$ and $\operatorname{Var}[R(t)]=\beta^{2}\,t^{2H}$. Hence, the $t^{H}$ law (1) arises theoretically:
@@
Observe that while $\mathbb{E}[R(t)]$ depends on the time-dependent drift term $\alpha(t)$, $\operatorname{SD}[R(t)]$ is determined solely by the Hurst parameter $H$ and the constant volatility $\beta$, remaining independent of $\alpha(t)$. Since $B_{H}(t)$ is Gaussian and $R(t)$ is a linear transformation of it (via Eq. (14)), $R(t)$ itself follows a Gaussian distribution. Thus, for any fixed $t$, $X(t)$ follows a log-normal distribution with mean $\ln X_{0}+\mathbb{E}[R(t)]$ and variance $\operatorname{Var}[R(t)]$.
@@ III.5 Interpolating between log-normal and power-law distributions
Such conditions include the system being anti-persistent, i.e. $H\in(0,\tfrac{1}{2})$, having a moderate $\beta$ value, and an $\alpha(t)$ that does not induce excessive fluctuations in the baseline trajectory of $X(t)$. The last condition is empirically satisfied, as a typical average citation history curve follows a jump-and-decay pattern.
To clarify, if we approximate $C(T)$ as $\sum_{i=1}^{n}X(t_{i})\mathop{}\!\Delta t$ (i.e. treating it as a sum of random variables) and $u(t)$ has a damping effect that causes $X(t)$ to rapidly approach zero, then for sufficiently large $i$, the terms $X(t_{i})\mathop{}\!\Delta t$ become negligible. As a result, each individual term’s contribution to the integral is suppressed, reducing the likelihood of heavy tails. Moreover, even when $X(t_{i})$ reaches a high value at a given instant, anti-persistence naturally limits the occurrence of consecutive large peaks. If $\beta$ is also moderate, the probability of extreme peaks decreases further. Thus, under these conditions, the distribution of $C(T)$ tends to form a single, relatively compact ‘bump’ with its peak near $C=0$ rather than an extended heavy tail.
@@ IV Numerical Simulations
Based on the theory we have developed, we perform numerical simulations of $X(t)$ and $C(T)$ to verify whether the theoretical predictions discussed in the previous section are reproduced and whether the observed macroscopic picture of citation dynamics is appropriately replicated. We primarily focus on the anti-persistent case, while the persistent case is addressed in the supplementary materials [53].
@@ IV.1 Simulation setup
For the numerical implementation of fractional Brownian motion, the fbm package [69] in Python was used. The Hurst parameter was set to $H=0.15$, ensuring a sufficiently strong anti-persistent nature of the system. The fractional SDE (8) was discretised with a time step of $n=1{,}000$ and a final time of $T=10$, yielding $\Delta t=T/n=10^{-2}$. As a toy model for the function representing the average attention history curve $u(t)$, the log-normal probability distribution function (PDF) was employed. To ensure that $X_{0}>0$, we used a rescaled version, given by
@@
Here, $\omega$ is the shape parameter, $e^{\theta}$ represents the median of the log-normal PDF, and $t_{0}>0$ is the location parameter.
@@ IV.2 Simulating attention curves
With the above implementation and parameter settings, the numerical simulation results for the trajectory of $\{\hat{X}_{i}\}_{i=1}^{n}$ are shown in Fig. 3. For $\beta=0.2$ (Fig. 3(a)), the trajectory closely follows a log-normal PDF, as theoretically expected. As $\beta$ increases ((a) to (b) to (c)), fluctuations become more pronounced, and the trajectory exhibits greater instability, aligning with theoretical predictions. Furthermore, due to the decay effect of $\hat{u}(t)$, the period during which the attention undergoes large fluctuations remains relatively short. Beyond a certain point, only negligible fluctuations persist, further reinforcing theoretical expectations.
Additionally, an increase in volatility raises the likelihood of rare but extreme values of $\hat{X}$; for example, note the high peak of $\hat{X}>300$ in Fig. 3(c). However, a crucial aspect of the current setup—where $H$ is significantly less than $0.5$, indicating a strongly anti-persistent system—is that such peaks do not persist or occur repeatedly over extended periods. Instead, they appear as relatively isolated and sporadic events. In other words, when $\hat{X}$ reaches a large value at a given moment, it is likely to decrease soon after due to a counteracting effect. Consequently, contributions to the integral become more dispersed, reducing the tendency for the distribution of $\hat{C}(T)$ to develop an excessively heavy tail.
Figure 2: Simulation results of the attention curve $\hat{X}(t)$ for an anti-persistent system ($H=0.15$). The parameters in $\hat{\alpha}(t)$ are set to $\theta=0.48$ and $\omega=0.8$. The volatility values are (a) $\beta=0.2$, (b) $\beta=0.8$ and (c) $\beta=1.5$.
@@ IV.3 Monte Carlo simulation of citation distributions
Under the same settings described above, we performed $N_{\mathrm{s}}=50{,}000$ Monte Carlo simulations to examine the distribution of cumulative citation counts up to time $T$. Here, $N_{\mathrm{s}}$ can be simply interpreted as the number of publications in the similated world. The simulated citation count was obtained as follows [53]:
@@
[…]
@@
Table 1: Summary of the estimated shape parameter ($\hat{\sigma}$) for the fitted log-normal distribution under varying volatility parameters ($\beta$), given $H=0.15$ (anti-persistent) and $\hat{\alpha}(t)$ from Eq. (16) with $\theta=0.48$ and $\omega=0.8$. All cases satisfy $p<0.001$.
@@
For comparison, it is instructive to conduct a hypothetical simulation in which the system exhibits a persistent memory effect and attention does not decay [53]. In this scenario, compared to the anti-persistent case, the trajectory of attention becomes smoother, with less pronounced jagged fluctuations (Fig. S2). The distribution of $\hat{C}(T)$ develops a fatter tail, and this tendency becomes more pronounced as $\beta$ increases (Fig. S2(a) to (b) to (c)). Moreover, the high-citation region can be well approximated by a shifted power law (Fig. S3(a)–(c)). All these findings align with the theoretical predictions presented in the previous section.
@@ IV.4 Simulating the $\bm{t^{H}}$ law
Finally, we confirm that the Monte Carlo simulations successfully replicate the $t^{H}$ law observed in Fig. 1. We used the same settings as in the earlier anti-persistent scenario, except for assigning $X_{0}=1$. For the ‘yearly’ time intervals in Fig. 1, we set $\tau_{k}=k\in\mathbb{N}$. Then, for each period $\{(\tau_{k-1},\tau_{k}]\}_{k=1}^{10}$, we compute the number of citations acquired during year $k$, denoted as $\hat{c}_{k}$, via the same method as Eq. (17):
@@
[…]
@@ V.1 Exploring citation dynamics in arXiv
We analysed the relationship between the number of citations received annually and the time elapsed since posting for arXiv e-prints published between 1995 and 2006, tracking citations up to February 2020. The dataset used is the same as that employed in [54, 23]. Since these citation data are counted annually, the number of data points is at most a few dozen, which is significantly smaller than the long-term datasets typically assembled in financial engineering [53]. Under these conditions, the $t^{H}$ law presented in Eq. (15) serves as a key methodological guide for estimating the fundamental parameters of the system. Leveraging this relation, we determine the values of $H$ and $\beta$ through a straightforward regression analysis.
@@
For each individual e-print in the dataset, we compiled citation data annually, defining the elapsed years as $\tau_{k}=k\in\mathbb{N}$ with $\Delta\tau=\tau_{k}-\tau_{k-1}=1$. Specifically, we constructed the dataset $\{\Delta C_{k}\}$, representing the number of citations received in year $k$, i.e., $k$ years since publication, and then computed $\big{\{}R_{k}\coloneqq\ln\big{(}\Delta\widetilde{C}_{k}/\Delta\tau\big{)}\big{\}}$ in accordance with Eq. (14). Here, we define $\Delta\widetilde{C}_{k}=\Delta C_{k}+\epsilon$ with $\epsilon=10^{-15}$, introducing a cut-off to prevent divergence when $\Delta C_{k}=0$. The bias introduced by this adjustment is assumed to be negligible [70]. Finally, for each year, we computed $\{\operatorname{SD}[R_{k}]\}$. Performing a linear regression on the obtained data points $\mathrm{P}_{k}\equiv(\ln\tau_{k},\ln\operatorname{SD}[R_{k}])$, we estimated the upper bound of $H$ from the slope. Similarly, the lower bound of $\beta$ was obtained by exponentiating the intercept of the regression model. It is important to note that $\operatorname{SD}[R_{k}]$ in this regression represents the SD of citation counts during year $k$, which is distinct from the SD of the logarithm of cumulative citations over $[0,T]$, denoted as $\sigma$ earlier.
Figure S4 presents scatter plots of $\{\mathrm{P}_{k}\}$ for each publication year, overlaid with regression lines, while Table S2 summarises the corresponding linear regression results. For all groups of e-prints published in different years, the $t^{H}$ law holds well, with the results for 2001 matching those in Fig. 1. The estimated values $\hat{H}$ and $\hat{\beta}$ over time are shown in Figs. S5(a) and (b), respectively. These remain approximately within the ranges of $0.08$ to $0.15$ for $\hat{H}$, with an average of $\hat{H}\approx 0.13$, and $0.53$ to $0.65$ for $\hat{\beta}$, without any noticeable deviations or trends contradicting constancy.
@@
While these results, derived from a simplified estimation, should be considered with the aforementioned biases in mind, they still provide statistically solid and qualitatively robust insights into memory effects in citation dynamics. From these estimations, we conclude that the citation dynamics of e-prints on arXiv follow a strongly anti-persistent stochastic process. Combining this with the simulation insights from Section IV strengthens the theoretical basis for why the cumulative citation distribution of arXiv e-prints (as observed in [23]) is well approximated by a highly skewed log-normal distribution with a pronounced peak near $C=0$.
@@ V.2 Hidden fractality in citation trajectory
Fractal phenomena in human social activities have been explored in various contexts, including urban morphology and city structures [71], financial markets [72], and complex networks [73, *Song06]. However, quantitatively estimating fractal dimensions in such systems remains challenging. In citation dynamics, although the above finding is indirect rather than based on direct observation, it is noteworthy that it provides quantitative insight into the underlying intrinsic fractal structure.

### v2

@@ III.4 The fractional stochastic differential equation
Now we explore the infinitesimal-time limit $\Delta t\to dt$. To incorporate our set of Ansätze (i)(iv), we consider the following fractional SDE, which describes the evolution of latent attention:
@@
The time-dependent drift function $\alpha(t)$ determines the shape of the average attention history curve $X(t)$ [see Eq. (11) below]. The parameter $\beta>0$, referred to as the volatility parameter—borrowing terminology from financial engineering—controls the magnitude of random fluctuations in $X(t)$ over the interval $dt$, reflecting external random events as discussed earlier.
As is evident, the fractional SDE (8) inherently incorporates our Ansätze (iii) and (iv) by definition. In particular, Ansatz (iii) satisfies the property of nonlinear preferential attachment. Compared to the SDE in the Black–Scholes model [51, *Merton73], this fractional SDE differs in two key aspects: Brownian motion is replaced by fractional Brownian motion, and the drift term $\alpha(t)$ is a function of time rather than a constant. If $\alpha$ were constant, then this model would coincide with the fractional Black–Scholes model [67]. In addition, compared to the SDE considered in Eqs. (14) and (S3) in Ref. [23], this fractional SDE differs in that Brownian motion is extended to fractional Brownian motion, while the volatility remains a time-independent constant.
Next we consider solving the fractional SDE (8). If $H=\tfrac{1}{2}$, then the process is a semimartingale, allowing the SDE to be solved using Itô integration (Appendix C in Ref. [23]). However, in the fractional SDE (8), the process is nonsemimartingale, meaning that the integral $\int_{0}^{t}\mathop{}\!dB_{H}(s)$ cannot be evaluated in the Itô sense. Nonetheless, this integral is well-defined under the fractional Wick–Itô integral [68]. The solution is:
@@
Using the properties of fractional Brownian motion given by Eqs. (4) and (5), we obtain $\mathbb{E}[R(t)]=A(t)$ and $\operatorname{Var}[R(t)]=\beta^{2}\,t^{2H}$. Hence, the $t^{H}$ law (1) arises theoretically:
@@
Observe that while $\mathbb{E}[R(t)]$ depends on the time-dependent drift term $\alpha(t)$, $\operatorname{SD}[R(t)]$ is determined solely by the Hurst parameter $H$ and the constant volatility $\beta$, remaining independent of $\alpha(t)$. Since $B_{H}(t)$ is Gaussian and $R(t)$ is a linear transformation of it [via Eq. (14)], $R(t)$ itself follows a Gaussian distribution. Thus, for any fixed $t$, $X(t)$ follows a log-normal distribution with mean $\ln X_{0}+\mathbb{E}[R(t)]$ and variance $\operatorname{Var}[R(t)]$.
@@ III.5 Interpolating between log-normal and power-law distributions
Such conditions include the system being antipersistent, i.e. $H\in(0,\tfrac{1}{2})$, having a moderate $\beta$ value, and an $\alpha(t)$ that does not induce excessive fluctuations in the baseline trajectory of $X(t)$. The last condition is empirically satisfied, as a typical average citation history curve follows a jump-and-decay pattern.
To clarify, if we approximate $C(T)$ as $\sum_{i=1}^{n}X(t_{i})\mathop{}\!\Delta t$ (i.e. treating it as a sum of random variables) and $u(t)$ has a damping effect that causes $X(t)$ to rapidly approach zero, then for sufficiently large $i$, the terms $X(t_{i})\mathop{}\!\Delta t$ become negligible. As a result, each individual term’s contribution to the integral is suppressed, reducing the likelihood of heavy tails. Moreover, even when $X(t_{i})$ reaches a high value at a given instant, antipersistence naturally limits the occurrence of consecutive large peaks. If $\beta$ is also moderate, then the probability of extreme peaks decreases further. Thus, under these conditions, the distribution of $C(T)$ tends to form a single, relatively compact ‘bump’ with its peak near $C=0$ rather than an extended heavy tail.
@@ IV Numerical Simulations
Based on the theory we have developed, we perform numerical simulations of $X(t)$ and $C(T)$ to verify whether the theoretical predictions discussed in the previous section are reproduced and whether the observed macroscopic picture of citation dynamics is appropriately replicated. We primarily focus on the antipersistent case, whereas the persistent case is discussed in the supplemental material [53]
@@ IV.1 Simulation setup
For the numerical implementation of fractional Brownian motion, the fbm package [70] in Python was used. The Hurst parameter was set to $H=0.15$, ensuring a sufficiently strong antipersistent nature of the system. The fractional SDE (8) was discretised with a time step of n=1,000n=1{,}000 and a final time of $T=10$, yielding $\Delta t=T/n=10^{-2}$. As a toy model for the function representing the average attention history curve $u(t)$, the log-normal probability distribution function (PDF) was employed. To ensure that $X_{0}>0$, we used a rescaled version, given by
@@
Here $\omega$ is the shape parameter, $e^{\theta}$ represents the median of the log-normal PDF, and $t_{0}>0$ is the location parameter.
@@ IV.2 Simulating attention curves
With the above implementation and parameter settings, the numerical simulation results for the trajectory of $\{\hat{X}_{i}\}_{i=1}^{n}$ are shown in Fig. 3. For $\beta=0.2$ [Fig. 3(a)], the trajectory closely follows a log-normal PDF, as theoretically expected. As $\beta$ increases [Figs. 3(a) to 3(b) to 3(c)], fluctuations become more pronounced, and the trajectory exhibits greater instability, aligning with theoretical predictions. Furthermore, due to the decay effect of $\hat{u}(t)$, the period during which the attention undergoes large fluctuations remains relatively short. Beyond a certain point, only negligible fluctuations persist, further reinforcing theoretical expectations.
Additionally, an increase in volatility raises the likelihood of rare but extreme values of $\hat{X}$; for example, note the high peak of $\hat{X}>300$ in Fig. 3(c). However, a crucial aspect of the current setup—where $H$ is significantly less than $0.5$, indicating a strongly antipersistent system—is that such peaks do not persist or occur repeatedly over extended periods. Instead, they appear as relatively isolated and sporadic events. In other words, when $\hat{X}$ reaches a large value at a given moment, it is likely to decrease soon after due to a counteracting effect. Consequently, contributions to the integral become more dispersed, reducing the tendency for the distribution of $\hat{C}(T)$ to develop an excessively heavy tail.
Figure 2: Simulation results of the attention curve $\hat{X}(t)$ for an antipersistent system ($H=0.15$). The parameters in $\hat{\alpha}(t)$ are set to $\theta=0.48$ and $\omega=0.8$. The volatility values are (a) $\beta=0.2$, (b) $\beta=0.8$ and (c) $\beta=1.5$.
@@ IV.3 Monte Carlo simulation of citation distributions
Under the same settings described above, we performed $N_{\mathrm{s}}=50{,}000$ Monte Carlo simulations to examine the distribution of cumulative citation counts up to time $T$. Here $N_{\mathrm{s}}$ can be simply interpreted as the number of publications in the similated world. The simulated citation count was obtained as follows [53]:
@@
where $\lceil\cdot\rceil$ denotes the ceiling function. The resulting histogram is shown in Fig. 3. As volatility increases [Fig. 3(a) to 3(b) to 3(c)], the reproduced distribution becomes closer to the highly skewed distributions commonly observed empirically. Moreover, given the present settings for $H$ and $\hat{\alpha}(t)$, $\hat{C}(T)$ tends not to develop a heavy tail for any choice of $\beta$, making it more amenable to approximation by a single log-normal distribution.
@@
Table 1: Summary of the estimated shape parameter ($\hat{\sigma}$) for the fitted log-normal distribution under varying volatility parameters ($\beta$), given $H=0.15$ (antipersistent) and $\hat{\alpha}(t)$ from Eq. (16) with $\theta=0.48$ and $\omega=0.8$. All cases satisfy $p<0.001$.
@@
For comparison, it is instructive to conduct a hypothetical simulation in which the system exhibits a persistent memory effect and attention does not decay [53]. In this scenario, compared to the antipersistent case, the trajectory of attention becomes smoother, with less pronounced jagged fluctuations (Fig. S2). The distribution of $\hat{C}(T)$ develops a fatter tail, and this tendency becomes more pronounced as $\beta$ increases [Fig. S2(a) to S2(b) to S2(c)]. Moreover, the high-citation region can be well approximated by a shifted power law [Fig. S3(a)–S3(c)]. All these findings align with the theoretical predictions presented in the previous section.
@@ IV.4 Simulating the $\bm{t^{H}}$ law
Finally, we confirm that the Monte Carlo simulations successfully replicate the $t^{H}$ law observed in Fig. 1. We used the same settings as in the earlier antipersistent scenario, except for assigning $X_{0}=1$. For the ‘yearly’ time intervals in Fig. 1, we set $\tau_{k}=k\in\mathbb{N}$. Then, for each period $\{(\tau_{k-1},\tau_{k}]\}_{k=1}^{10}$, we compute the number of citations acquired during year $k$, denoted as $\hat{c}_{k}$, via the same method as Eq. (17):
@@
Figure 5: Scatter plot of $(\tau_{k},\hat{c}_{k})$ (filled markers) and $(\tau_{k},\hat{X}(\tau_{k}))$ (unfilled markers) from the same data as in Fig. 3, for $k=1\text{--}10$, along with their linear regression fits. In both cases, the error bars are negligibly small and are therefore omitted for clarity.
@@ V.1 Exploring citation dynamics in arXiv
We analysed the relationship between the number of citations received annually and the time elapsed since posting for arXiv e-prints published between 1995 and 2006, tracking citations up to February 2020. The dataset used is the same as that employed in Refs. [55, 23]. Note that the arXiv dataset includes papers that were subsequently published in peer-reviewed journals, and that the citation data obtained from Semantic Scholar [71] include citations both from and to peer-reviewed publications.
Since these citation data are aggregated on an annual basis, the number of available data points is at most a few dozen, which is substantially smaller than the long-term datasets typically employed in financial engineering [53]. Because of this limitation, we cannot effectively apply widely used techniques such as rescaled range analysis (R/S) [72, 73] or detrended fluctuation analysis (DFA) [74, *Hu01, *Chen02]. Under such constraints, the $t^{H}$ law presented in Eq. (15) serves as a key methodological guideline for estimating the fundamental parameters of the system. By leveraging this relation, we determine the values of $H$ and $\beta$ through a straightforward regression analysis.
@@
For each individual e-print in the dataset, we compiled citation data annually, defining the elapsed years as $\tau_{k}=k\in\mathbb{N}$ with $\Delta\tau=\tau_{k}-\tau_{k-1}=1$. Specifically, we constructed the dataset $\{\Delta C_{k}\}$, representing the number of citations received in year $k$, i.e., $k$ years since publication, and then computed $\big\{R_{k}\coloneqq\ln\big(\Delta C_{k}/\Delta\tau\big)\big\}$ in accordance with Eq. (14). Here citation records with $\Delta C_{k}=0$ were excluded from the analysis. Finally, for each year, we computed $\{\operatorname{SD}[R_{k}]\}$. Performing a linear regression on the obtained data points $\mathrm{P}_{k}\equiv(\ln\tau_{k},\ln\operatorname{SD}[R_{k}])$, we estimated the upper bound of $H$ from the slope. Similarly, the lower bound of $\beta$ was obtained by exponentiating the intercept of the regression model. It is important to note that $\operatorname{SD}[R_{k}]$ in this regression represents the SD of citation counts during year $k$, which is distinct from the SD of the logarithm of cumulative citations over $[0,T]$, denoted as $\sigma$ earlier.
Figure S4 presents scatter plots of $\{\mathrm{P}_{k}\}$ for each publication year, overlaid with regression lines, while Table S2 summarises the corresponding linear regression results. For all groups of e-prints published in different years, the $t^{H}$ law holds well, with the results for 2001 matching those in Fig. 1. The estimated values $\hat{H}$ and $\hat{\beta}$ over time are shown in Figs. S5(a) and S5(b), respectively. These remain approximately within the ranges of $0.08$ to $0.15$ for $\hat{H}$, with an average of $\hat{H}\approx 0.13$, and $0.53$ to $0.65$ for $\hat{\beta}$, without any noticeable deviations or trends contradicting constancy.
The observed $t^{H}$ law can be interpreted as follows. Most papers receive no citations from the outset or eventually fall to $\Delta C_{k}=0$, forming the sharp peak at zero in the distribution. A small subset of long-lived papers maintain $\Delta C_{k}>0$, exhibiting diverse citation trajectories. Consequently, the distribution becomes a mixture: a largely static mass at zero and a gradually widening tail of nonzero values. The $t^{H}$ law reflects the increasing heterogeneity in $\ln(\Delta C_{k})$ within this tail over time.
@@
It should be noted that these results are obtained through a simplified estimation procedure and based on data with a limited number of points, and therefore must be interpreted with the aforementioned biases in mind. Nevertheless, they offer statistically solid and qualitatively robust insights into memory effects in citation dynamics. From these estimations, we conclude that the citation dynamics of e-prints on arXiv follow a strongly antipersistent stochastic process. Combining this with the simulation insights from Sec. IV strengthens the theoretical basis for why the cumulative citation distribution of arXiv e-prints (as observed in Ref. [23]) is well approximated by a highly skewed log-normal distribution with a pronounced peak near zero citations.
@@ V.2 Hidden fractality in citation trajectory
Fractal phenomena in human social activities have been explored in various contexts, including urban morphology and city structures [77], financial markets [73], and complex networks [78, *Song06]. However, quantitatively estimating fractal dimensions in such systems remains challenging. In citation dynamics, although the above finding is indirect rather than based on direct observation, it is noteworthy that it provides quantitative insight into the underlying intrinsic fractal structure.